These are the only two problems I was totally stuck on:
note: the 3's are like log10base
1.) Express the expression LOG3a+2log3b-log3c as a single logarithm expression.
2.) Express LOGb12 in terms of LOGb2 and LOGb3.

I don't even know how to start

Answer by stanbon(48535)

You have to know how to use the rules of logarithms.
1)I'm not going to type all the "a's" as you know it
is all base a:
You get log[(ab^2)/c]
2)Here I will not type the base "b".
log 12 = log (4*3) = log 4 + log 3 = log 2^2 + log 3 = 2log2 + log 3
If you can understand this you understand two of the log laws.
Cheers,
Stan H.